OpenStudy (anonymous):

if $300 is deposited in an account and it increases to $450 in eight years compounded quarterly. find the rate of interest to the nearest tenth

OpenStudy (anonymous):

A(t) = S * (1+P)^Y S= starting amount A = current Amount P = rate Y = amount of times Y =32 A(t) = 450 S = 300 Therefore P = (A(t)/S)^1/Y - 1 = 0.0127513991 or 1.275%

OpenStudy (anonymous):

that's per quarter btw :D For year: 5.10 %